OpenStudy (anonymous):

apply the quadratic formula to find the roots of the given function, and then graph the function. g(x) = 5x^2+13x-6 please show work for credit

OpenStudy (anonymous):

so you want it all written out? use \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=5,b=13,c=-6\] you get \[x=\frac{-13\pm\sqrt{13^2-4\times 5\times -6}}{2\times 5}\] \[x=\frac{-13\pm\sqrt{169+120}}{10}\]

OpenStudy (anonymous):

yes please I am having a hard time understanding this

OpenStudy (anonymous):

then we simplify right

OpenStudy (anonymous):

what happened to my post? i wrote it all out

OpenStudy (anonymous):

ii guess all that is left is \[x=\frac{-13\pm\sqrt{289}}{10}\]

OpenStudy (anonymous):

I don't know you disappered from the screen

OpenStudy (anonymous):

something funny going on here. ok did you get \[x=\frac{-13\pm\sqrt{289}}{10}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now \[\sqrt{289}=17\] so your two answers are \[\frac{-13-17}{10}=\frac{-30}{10}=-3\]

OpenStudy (anonymous):

and \[\frac{-13+17}{10}=\frac{4}{10}=\frac{2}{5}\]

OpenStudy (anonymous):

hope these went through ok. two solutions, -3 and 2/5

OpenStudy (anonymous):

so now what do I graph?